# The Narrow Depth and Breadth of Corporate Responsible AI Research

### v1

@@ 5.4 Frequency count
Additionally, we searched the abstracts of the same 10,799 papers used in our prior k-means clustering analysis. We used regular expressions (regex) to identify relevant papers and counted the total number of papers on each topic from industry and academia. To get an intuitive measure of industry’s focus on different topics, we calculated the percentage of papers on each topic by dividing the total number of papers from industry on each one (identified using regex patterns) by the total number of papers from industry on responsible AI (classified by our earlier model described in Section 4.1.2).
Repeating the same calculations for academia, and comparing the results with those of industry, we observed their contrasting focus on different topics. We found that industry has limited focus on issues like “human rights,” “environment concerns,” and “beneficence” (Fig. 3b) while paying more attention to “explainable AI.” This disparity suggests that industry may be prioritizing the development of AI systems that are easier to bring to market, while largely neglecting the potential negative externalities associated with AI. For example, the increasing computational power required for AI systems can contribute to significant environmental concerns, such as high energy consumption and carbon emissions, yet industry seems to be paying very limited attention to these issues. Overall, while both industry and academia converse somewhat similarly on issues like “non-maleficence,” on critical issues related to the negative implications of AI, they have a relatively divergent focus.
@@ 5.5 Additional Linguistic Analyses
Furthermore, we employed complementary approaches on a subset of the dataset previously analyzed, specifically focusing on papers from three leading conferences in the field of responsible AI. First, we used latent Dirichlet allocation (LDA) topic modeling (Blei et al.,, 2001). The results mirrored our prior findings from STM analysis as previously shown in Fig. 3b. The patterns held even when we varied the number of topics. In Fig. S10 in the appendix, we present LDA topic modeling results for ten topics, and we list the top keywords in Table LABEL:table:oldA8 in the appendix.
Finally, we used a bigram frequency analysis, which requires fewer assumptions than topic modeling and clustering approaches. Fig. S14 in the appendix shows the top 20 keywords for industry-only and academia-only bigrams using a different dataset containing the abstracts of papers from the three leading responsible AI conferences. This analysis suggests that industry research has limited engagement with terms like “criminal justice,” “implicit bias,” and “fairness metrics.” This alternative method reaffirmed our earlier findings that industry research tends to focus less on ethical concerns and societal impacts, favoring a greater emphasis on technical progress. Industry’s responsible AI research tends to focus on areas that align with a market-driven approach, such as "explainable AI" and "model performance." Firms have greater incentives to develop and launch products that are explainable, as this can help satisfy regulatory requirements and gain market acceptance.
@@ 5.6 Summary of Findings
Our linguistic analysis, employing diverse methods across different datasets and varying prior assumptions, revealed the distinct research priorities of industry and academia. The findings indicate that industry places less emphasis on key aspects of responsible AI compared to academia, instead focusing more on the technical aspects of responsible AI and areas that align with market demands. In contrast, academia pays more attention to crucial issues such as human rights considerations, moral concerns, and the societal implications of AI, which is more consistent with the “sociotechnical approach” needed to mitigate the impacts of advanced AI (Lazar and Nelson,, 2023; Birhane et al.,, 2022). Overall, it seems that industry-led AI research is on a trajectory that prioritizes solutions relying solely on technical means, without incorporating broader societal insights, which may compound AI’s dangers rather than address them effectively.
@@ 6 Limited Adoption of Responsible AI Research in Commercialization: Patent Citation Analysis
To assess the extent to which industry integrates responsible AI research into its product commercialization processes, we employed patent citations as a measure of how industry builds upon academic research in its products and services, following established practices in prior literature (Cao et al.,, 2023; Trajtenberg and Jaffe,, 2002). The intuition is that if AI firms incorporate responsible AI research in their products, it will likely be reflected in their patent citations (Trajtenberg and Jaffe,, 2002). This can be seen in the example provided in Fig. 4, which illustrates how a patent citation can reveal the impact of responsible AI research on firms’ innovations.
@@ 6.1 Data and Methodology:
Prior research suggests that industry patents tend to pay more attention to industry research and cite their own research more in their commercial innovations (Bikard,, 2018; Arora et al.,, 2021). To consider this different attention pattern, we conducted our patent citation analysis for industry and academia co-authored papers separately. We compiled two new, distinct datasets of responsible AI papers, separating papers from industry and academia.
@@ 6.2 Responsible AI Research in Generic Patents
Our analysis of over 32 million patent citations in USPTO patents shows that industry patents rarely cite responsible AI research. We found that only 88 industry-authored responsible AI papers were cited in these generic patents from 2010 to 2022, as presented in Fig. 5a. In contrast, generic patents cited 7,532 conventional AI papers and 26,970 non-AI papers from industry over the same period. [^9] A notable trend is the decreasing total number of citations across both categories–conventional AI and responsible AI (see Fig. S20 in the appendix for the cumulative counts). There could be two potential explanations for this trend. First, it takes time for basic knowledge to diffuse to commercial patents (Mansfield,, 1991; Cao et al.,, 2023). Also, older papers have more time to be discovered by others. Therefore, patents are likely to cite older papers more than newer papers. The second reason is the time-consuming patent application process; our data considers only the approved patents, not patent applications still under review, contributing to the reduced citation count in recent years. We found a total of 246 patents cited 88 responsible AI papers. Then we examined the top patent assignees’ that cited industry-authored responsible AI research: IBM, Microsoft, and Google were the top patent assignees. Interestingly, these are the same set of firms that had more engagement in responsible AI research. The correlation between firms’ engagement in responsible AI research and their patent citations emphasizes the crucial role of absorptive capacity (Cohen and Levinthal,, 1990) in facilitating the translation of responsible AI research into commercial inventions.
We observed a similar pattern when we examined industry patents’ citations of academic papers. By matching over 5 million academic papers from Scopus (2010-22) with data on over 32 million patent-paper-citations (1947-2022), we observed that industry cited 15,236 conventional AI and 140,508 non-AI academic papers, but only 396 responsible AI papers from academia (see Fig. 5c). Upon examining the patents’ citations we found a total of 983 patents cited academic responsible AI papers. Once again, the top patent assignees citing such academic papers were IBM, Microsoft, and Google. [^10] When we manually checked these responsible AI papers we found a slight over-count of responsible AI papers in industry and academia because of the keywords sampled responsible AI papers.
@@
Figure 5: This figure analyzes industry and academic papers’ citations in USPTO patents. By matching against a comprehensive list of USPTO patent-paper citation data (n = 32,698,465 citations; 1947-2022), we show in Fig. 5a and 5c that, only 88 and 396 responsible AI papers from industry and academia, respectively, have been cited in generic patents, while 15,236 academia and 7,532 industry conventional AI papers have been cited in patents between 2010-22. Using a separate dataset of AI patents (n = 141,770 patents; 1985-2018), Fig. 5b and 5d illustrate that three responsible AI papers from industry and 17 from academia have been cited in the AI patents between 2010-18, in stark contrast to the extensive incorporation of conventional AI research.
@@ 6.3 Responsible AI Research in AI Patents
We further examined a subset of patents–specifically AI patents using a different source. Our result, presented in Fig. 5b shows that out of 141,770 AI patents (1985-2018), cited only three responsible AI papers from industry. However, during the same period, the patents cited 1,438 conventional AI papers and 908 non-AI papers.
Similarly, we analyze academic papers cited in AI patents from industry. Our prior pattern was also found here–only 17 responsible AI paper from academia were cited in AI patents (see Fig. 5d). In contrast, 1,949 conventional AI and 2,541 non-AI papers were cited by those patents. This disparity implies that while industry may draw upon academic conventional AI research, it does not seem to do the same for academic responsible AI research. However, in Fig. S19, we present evidence that AI firms’ responsible AI research builds on academic responsible AI research.
We then analyze the patents that cite the responsible AI papers from industry and academia. We find that the 17 university-authored responsible AI papers were cited by a total of 22 patents and 3 industry-authored responsible AI papers received citations from 5 patents.
@@ 6.4 Summary of Findings
Taken together, our large-scale patent-to-paper citation analysis is consistent with our prior results. It shows that industry has limited engagement in responsible AI, suggesting that firms are barely integrating responsible AI research into their commercial products. This limited integration could stem from two potential factors: either firms are not dedicating adequate attention to responsible AI research (Bikard,, 2018), or they lack the absorptive capacity (Cohen and Levinthal,, 1990; Zahra and George,, 2002) necessary to effectively assimilate and apply the knowledge produced by such research.
@@ 7 Discussion
Our results suggest that the rapid commercialization of AI (Maslej et al.,, 2023) is taking place without sufficient consideration of its potential consequences. Industry’s limited engagement with responsible AI research is concerning because AI firms are currently at the forefront of research and innovation, shaping the trajectory of AI development. Yet our results show that they have a limited engagement with responsible AI in both research (as measured by publications) and development (as measured by patent citations of responsible AI research). This means that the current AI trajectory is on a path that might not be socially optimal.
@@
This lack of engagement in research is also a cause for concern because innovation research suggests that firms need to participate in research to develop their capacity to absorb external knowledge (Cohen and Levinthal,, 1990; Zahra and George,, 2002). AI firms, however, are starting to reduce their responsible AI research teams (Vynck and Oremus,, 2023). For instance, leading technology firms have fired their responsible AI researchers at a disproportionate rate (Field and Vanian,, 2023) (see Table LABEL:table:AI-layoffs for a list of recent firings). [^11] Prior research further reveals that employees who prioritize ethical considerations often encounter obstacles, as firms tend to prioritize product launches over investing resources in addressing ethical concerns (Ali et al.,, 2023). If industry does not engage in responsible AI research, its ability to adopt the latest external research will be limited. Consequently, this lack of engagement could lead to less critical examinations of social benefits and costs in the early stages of the design and could inhibit the diffusion of responsible AI ideas from academia to industry.
While it is crucial for industry to enhance its engagement with the public, it is essential to recognize that such involvement should supplement, not replace, public responsible AI research. There is a significant concern that industry might have excessive influence over the trajectory of AI research (Young et al.,, 2022; Ahmed and Wahed,, 2020; Benkler,, 2019) and AI firms might unduly influence the future research agenda in the field of responsible AI (Hao,, 2021; Young et al.,, 2022; Baker and Hanna,, 2022; Jazwinska,, 2022). Therefore, we argue that industry’s increased presence should be complementary to extant responsible AI research.
@@
Additionally, the responsible AI research community has a key role to play in making its work more accessible and applicable for practitioners. For example, prior studies (Holstein et al.,, 2019; Ali et al.,, 2023; Ghosh et al.,, 2023) have noted the difficulties practitioners face when attempting to apply and implement research findings in real-world settings. Moreover, a substantial portion of academic responsible AI research may not be immediately applicable or easily integrated into commercial products, potentially contributing to the limited industry engagement in this area. More efforts are needed to ensure that industry and academia work together to translate academic responsible AI research into commercial innovation.
@@ 7.1 Policy Implications
Our systematic analysis reveals a concerning lack of industry involvement in responsible AI research, which poses significant risks to society. The development and commercialization of AI technologies without adequate consideration of their societal implications could have far-reaching negative consequences. Therefore, it is important to implement comprehensive measures to ensure the responsible development and deployment of AI systems.
@@
Our results suggest that the lack of meaningful participation in responsible AI research is a key reason why firms are largely reactive to AI-induced challenges. This lack of meaningful participation hinders their ability to anticipate and mitigate the potential downsides of their tools. For example, in 2016, Microsoft launched an AI chatbot on Twitter targeting 18-to-24-year-olds, which within 24 hours of its launch started posting anti-semitic, racist, and misogynist content (Mason,, 2016). Years later, in 2023, Snapchat launched its AI chatbot, My AI, which reportedly encouraged physical intimacy with a minor (Fowler,, 2023). These examples, spanning several years, underscore the pervasive reactive stance of the AI ecosystem, where mitigating harms is often treated as an afterthought rather than a proactive design principle. By actively engaging in responsible AI research, firms can develop a more comprehensive understanding of the potential risks and ethical implications of their technologies, enabling them to proactively address these issues during the design and development process.
Despite the mounting pressure (The White House, 2023a, ; Krafft et al.,, 2021; Ahmed and Jia,, 2023; Belfield,, 2020) stemming from the rapidly rising number of incidents concerning the misuse of AI (Maslej et al.,, 2024), industry participation in responsible AI research remains limited. This suggests that existing incentives might not be sufficient for the corporate development of responsible AI (Askell et al.,, 2019). Prior research contends that market pressure could accelerate the commercialization process, leading firms to underinvest in responsible AI development (Askell et al.,, 2019). Regulatory frameworks could induce AI firms to demonstrate substantial engagement in responsible AI research as a prerequisite for product development and deployment.
@@
Transparency and public accountability are vital in the AI development process (Brundage et al.,, 2020; Mittelstadt,, 2019). To foster trust and confidence in the AI ecosystem, policymakers should mandate AI firms to disclose their responsible AI research efforts, ethical frameworks, and decision-making processes related to the development and deployment of AI systems (de Laat,, 2021; de Almeida et al.,, 2021). These disclosures should be subject to independent audits and public scrutiny, ensuring that AI development remains aligned with societal values and interests (Mökander et al.,, 2021).
@@
Furthermore, policymakers need to prioritize the allocation of funding and resources toward independent, public-sector responsible AI research initiatives. While industry participation is vital, our findings underscore the importance of maintaining a robust, impartial research ecosystem that can critically examine AI technologies and their societal impacts without being influenced by commercial interests (Baker and Hanna,, 2022).
@@
More industry-academia research collaborations in responsible AI could help to bridge the gap. These partnerships would promote knowledge exchange, foster the integration of academic insights into industry practices, and cultivate a culture of responsible AI development within firms. By leveraging the combined expertise and resources of academia and industry, such collaborations could drive the advancement of responsible AI research and its practical implementation.
7.2 Direction of AI is on a Suboptimal Path

### v2

@@ 5.4 Frequency count
Additionally, we searched the abstracts of the same 10,799 papers used in our prior k-means clustering analysis. We used regular expressions (regex) to identify relevant papers and counted the total number of papers on each topic from industry and academia. To get an intuitive measure of industry’s topic distribution, we calculated the percentage of papers on each topic by dividing the total number of papers from industry on each one (identified using regex patterns) by the total number of papers from industry on responsible AI (classified by our earlier model described in Section 4.1.2).
Repeating the same calculations for academia, and comparing the results with those of industry, we observed differences in topic distribution. We found that industry research shows lower frequencies for terms related to human rights,” environmental concerns,” and beneficence” (Fig. 3c) and higher frequencies for explainable AI.” This pattern is consistent with theoretical accounts suggesting that market-facing organizations may concentrate research efforts on areas with clearer paths to commercialization (Rosenberg and others, 1990; Cohen and Levinthal, 1990). For example, the increasing computational power required for AI systems can contribute to environmental concerns, such as high energy consumption and carbon emissions, a topic that appears less frequently in industry research than in academic research. Overall, while both industry and academia show similar frequencies for terms related to “non-maleficence,” their topic distributions differ for other dimensions of responsible AI research.
@@ 5.5 Additional Linguistic Analyses
Furthermore, we employed complementary approaches on a subset of the dataset previously analyzed, specifically focusing on papers from three leading conferences in the field of responsible AI. First, we used latent Dirichlet allocation (LDA) topic modeling (Blei, Ng, and Jordan, 2001). The results mirrored our prior findings from STM analysis as previously shown in Fig. 3b. The patterns held even when we varied the number of topics. In Fig. S11 in the appendix, we present LDA topic modeling results for ten topics, and we list the top keywords in Table LABEL:table:oldA8 in the appendix.
Finally, we used a bigram frequency analysis, which requires fewer assumptions than topic modeling and clustering approaches. Fig. S15 in the appendix shows the top 20 keywords for industry-only and academia-only bigrams using a different dataset containing the abstracts of papers from the three leading responsible AI conferences. This analysis suggests that industry research shows lower frequencies for terms like criminal justice,” implicit bias,” and fairness metrics.” These results are consistent with our earlier findings indicating that industry and academia differ in their topic distributions, with industry research more concentrated in technical topics. Industry’s responsible AI research shows higher frequencies in areas such as explainable AI” and “model performance,” which align with market-oriented applications. Firms may have incentives to develop and launch products that are explainable, as this can help satisfy regulatory requirements (Ahmed, Polidoro Jr, and Das, 2026) and gain market acceptance.
@@ 5.6 Summary of Findings
Our linguistic analysis, employing multiple methods across different datasets, documents variation in the thematic composition of responsible AI research between industry and academia. Industry research concentrates in technical aspects of responsible AI, such as explainability, privacy, and algorithmic fairness, areas that align closely with product development and market applications. Academia, by contrast, places greater emphasis on human rights considerations, ethical and moral dimensions, and the broader societal implications of AI, consistent with the sociotechnical approach discussed in recent literature (Lazar and Nelson, 2023; Birhane, Kalluri, Card, Agnew, Dotan, and Bao, 2022). These patterns hold across STM topic modeling, k-means clustering, frequency analysis, and LDA, suggesting they are robust to methodological choices. Taken together, the evidence indicates that industry and academia exhibit distinct thematic priorities within the responsible AI research landscape.
@@ 6 Limited Adoption of Responsible AI Research in Commercialization: Patent Citation Analysis
To assess the extent to which industry integrates responsible AI research into its product commercialization processes, we employed patent citations as a measure of how industry builds upon academic research in its products and services, following established practices in prior literature (Cao, Lu, Deng, Mcfarland, and Bernstein, 2023; Trajtenberg and Jaffe, 2002). The intuition is that if AI firms incorporate responsible AI research in their products, it will likely be reflected in their patent citations (Trajtenberg and Jaffe, 2002). This can be seen in the example provided in Fig. 4, which illustrates how a patent citation can reveal the impact of responsible AI research on firms’ innovations.
@@ 6.1 Data and Methodology:
Prior research suggests that industry patents tend to cite industry research at higher rates and cite their own research more in their commercial innovations (Bikard, 2018; Arora, Belenzon, and Sheer, 2021). To account for this citation pattern, we conducted our patent citation analysis for industry and academia co-authored papers separately. We compiled two new, distinct datasets of responsible AI papers, separating papers from industry and academia
@@ 6.2 Responsible AI Research in Generic Patents
Our analysis of over 32 million patent citations in USPTO patents shows that responsible AI research represents a small share of industry patent citations. We found that 88 industry-authored responsible AI papers were cited in these generic patents from 2010 to 2022, as presented in Fig. 5a. By comparison, generic patents cited 7,532 conventional AI papers and 26,970 non-AI papers from industry over the same period. [^9] A notable trend is the decreasing total number of citations across both categories–conventional AI and responsible AI (see Fig. S21 in the appendix for the cumulative counts). There could be two potential explanations for this trend. First, it takes time for basic knowledge to diffuse to commercial patents (Mansfield, 1991; Cao, Lu, Deng, Mcfarland, and Bernstein, 2023). Also, older papers have more time to be discovered by others. Therefore, patents are likely to cite older papers more than newer papers. The second reason is the time-consuming patent application process; our data considers only the approved patents, not patent applications still under review, contributing to the reduced citation count in recent years. We found a total of 246 patents cited 88 responsible AI papers. We then examined the top patent assignees that cited industry-authored responsible AI research: IBM, Microsoft, and Google were the top patent assignees. These are the same firms that had higher rates of responsible AI research output. The correlation between firms’ responsible AI research output and their patent citations is consistent with absorptive capacity theory (Cohen and Levinthal, 1990), which suggests that research participation facilitates the translation of knowledge into commercial inventions.
We observed a similar pattern when we examined industry patents’ citations of academic papers. By matching over 5 million academic papers from Scopus (2010-22) with data on over 32 million patent-paper-citations (1947-2022), we observed that industry cited 15,236 conventional AI and 140,508 non-AI academic papers, and 396 responsible AI papers from academia (see Fig. 5c). Upon examining the patents’ citations we found a total of 983 patents cited academic responsible AI papers. The top patent assignees citing such academic papers were IBM, Microsoft, and Google. [^10] When we manually checked these responsible AI papers we found a slight over-count of responsible AI papers in industry and academia because of the keywords sampled responsible AI papers.
@@
Figure 5: This figure analyzes industry and academic papers’ citations in USPTO patents. By matching against a comprehensive list of USPTO patent-paper citation data (n = 32,698,465 citations; 1947-2022), we show in Fig. 5a and 5c that 88 and 396 responsible AI papers from industry and academia, respectively, have been cited in generic patents, while 15,236 academia and 7,532 industry conventional AI papers have been cited in patents between 2010-22. Using a separate dataset of AI patents (n = 141,770 patents; 1985-2018), Fig. 5b and 5d illustrate that three responsible AI papers from industry and 17 from academia have been cited in the AI patents between 2010-18, compared to higher citation counts for conventional AI research.
@@ 6.3 Responsible AI Research in AI Patents
We further examined a subset of patents–specifically AI patents using a different source. Our result, presented in Fig. 5b shows that out of 141,770 AI patents (1985-2018), three responsible AI papers from industry were cited. During the same period, the patents cited 1,438 conventional AI papers and 908 non-AI papers.
Similarly, we analyzed academic papers cited in AI patents from industry. This analysis yielded comparable results–17 responsible AI papers from academia were cited in AI patents (see Fig. 5d). By comparison, 1,949 conventional AI and 2,541 non-AI papers were cited by those patents. This pattern indicates that citation rates for responsible AI research differ from those for conventional AI research in patent filings. In Fig. S20, we present evidence that AI firms’ responsible AI research builds on academic responsible AI research.
We then analyzed the patents that cite the responsible AI papers from industry and academia. We find that the 17 university-authored responsible AI papers were cited by a total of 22 patents and 3 industry-authored responsible AI papers received citations from 5 patents.
6.4 Alternative Channels for Responsible AI Integration
Patent citations represent one channel through which responsible AI research may inform commercial development, though firms may integrate responsible AI principles through additional mechanisms not captured by bibliometric analysis. Several firms have implemented responsible AI frameworks through governance structures, open-source tools, and certification processes. For example, IBM developed and open-sourced the AI Fairness 360 toolkit, which provides over 70 fairness metrics and bias mitigation algorithms for practitioners (Bellamy, Dey, Hind, Hoffman, Houde, Kannan, Lohia, Martino, Mehta, Mojsilović, and others, 2019). Salesforce established its Office of Ethical and Humane Use in 2018, developed an AI Ethics Maturity Model to guide organizational practices, and implemented consequence scanning methodologies across product teams (Salesforce, 2023). SAP, the first European technology firm to publish AI ethics principles, has achieved ISO 42001 certification for its AI governance framework covering Joule and SAP AI Core, and maintains both an internal Global AI Ethics Steering Committee and an external Global AI Ethics Advisory Panel (SAP, 2025). These examples illustrate that some firms operationalize responsible AI principles through product design, governance frameworks, and certification processes, activities that may not be reflected in traditional bibliometric measures of research output or patent citation patterns.
@@ 6.5 Summary of Findings
Taken together, our large-scale patent-to-paper citation analysis is consistent with our prior results. The analysis documents that industry patents cite responsible AI research at lower rates than conventional AI research, suggesting differences in how these research streams are integrated into commercial innovation. This pattern could reflect several factors: variation in firms’ research priorities, differences in the applicability of responsible AI findings to patentable innovations, or variation in absorptive capacity (Cohen and Levinthal, 1990; Zahra and George, 2002) necessary to assimilate and apply knowledge from this emerging field. Notably, firms with greater engagement in responsible AI research, such as IBM, Microsoft, and Google, are also the most frequent citers of responsible AI papers in their patents, consistent with the view that research engagement facilitates knowledge integration.
@@ 7 Discussion
Our results indicate that the rapid commercialization of AI (Maslej, Fattorini, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Ngo, Niebles, Parli, Shoham, Wald, Clark, and Perrault, 2023) is occurring concurrently with comparatively limited industry engagement in responsible AI research. This pattern has implications for the trajectory of AI development, as AI firms are currently at the forefront of research and innovation, shaping the field’s direction. Our findings document that firms exhibit lower engagement with responsible AI in both research (as measured by publications) and development (as measured by patent citations of responsible AI research). These patterns raise questions about whether the current trajectory of AI development adequately incorporates considerations of societal implications.
@@
These patterns of engagement also have implications for firms’ absorptive capacity. Innovation research suggests that firms benefit from participating in research to develop their capacity to absorb external knowledge (Cohen and Levinthal, 1990; Zahra and George, 2002). Recent reports indicate that some AI firms have reduced their responsible AI research teams (Vynck and Oremus, 2023), with several leading technology firms restructuring positions in this area. (Field and Vanian, 2023)  [^11] Prior research further reveals that employees who prioritize ethical considerations often encounter obstacles, as firms tend to prioritize product launches over investing resources in addressing ethical concerns (Ali, Christin, Smart, and Katila, 2023). If industry participation in responsible AI research remains limited, firms’ ability to adopt the latest external research may be constrained. This dynamic could affect the extent of critical examination of social benefits and costs in the early stages of design and may influence the diffusion of responsible AI knowledge from academia to industry.
While industry engagement with the public represents one avenue for advancing responsible AI, scholars have noted that such involvement functions most effectively when it supplements, rather than replaces, public responsible AI research. Prior work has raised questions about the potential for industry presence to shape the trajectory of AI research (Young, Katell, and Krafft, 2022; Ahmed and Wahed, 2020; Benkler, 2019) and influence the future research agenda in the field of responsible AI (Hao, 2021; Young, Katell, and Krafft, 2022; Baker and Hanna, 2022; Jazwinska, 2022). These considerations suggest that industry’s increased presence may be most beneficial when complementary to existing responsible AI research efforts.
@@
Additionally, the responsible AI research community plays an important role in making its work more accessible and applicable for practitioners. Prior studies (Holstein, Wortman Vaughan, Daumé, Dudik, and Wallach, 2019; Ali, Christin, Smart, and Katila, 2023; Ghosh, Kvitca, and Wilson, 2023) have documented challenges practitioners face when attempting to apply and implement research findings in real-world settings. Moreover, a substantial portion of academic responsible AI research may not be immediately applicable or easily integrated into commercial products, potentially contributing to the observed patterns of industry engagement in this area. Strengthening mechanisms for industry-academia collaboration could facilitate the translation of academic responsible AI research into commercial innovation.
@@ 7.1 Policy Implications
Our systematic analysis documents patterns of industry involvement in responsible AI research that have implications for AI governance. The development and commercialization of AI technologies intersects with questions about societal implications, which policymakers and scholars have increasingly sought to address. These findings may inform discussions about measures to support responsible development and deployment of AI systems.
@@
Our results suggest that patterns of participation in responsible AI research may be associated with firms’ capacity to anticipate challenges arising from AI deployment. Limited research engagement may constrain firms’ ability to identify and address potential issues with their tools during the design phase. Documented incidents illustrate this dynamic: in 2016, Microsoft launched an AI chatbot on Twitter targeting 18-to-24-year-olds, which within 24 hours of its launch began posting harmful content (Mason, 2016). Years later, in 2023, Snapchat launched its AI chatbot, My AI, which reportedly generated inappropriate content involving a minor (Fowler, 2023). These cases, spanning several years, illustrate how firms may encounter challenges that could potentially have been anticipated through greater engagement with responsible AI research. Active engagement in responsible AI research may enable firms to develop a more comprehensive understanding of the potential risks and ethical implications of their technologies, potentially allowing them to address these considerations during the design and development process.
Notwithstanding increased attention from regulators (The White House, 2023a), civil society (Krafft, Young, Katell, Lee, Narayan, Epstein, Dailey, Herman, Tam, Guetler, Bintz, Raz, Jobe, Putz, Robick, and Barghouti, 2021), and employees (Ahmed and Jia, 2023; Belfield, 2020), alongside a growing number of documented AI incidents (Maslej, Fattorini, Perrault, Parli, Reuel, Brynjolfsson, Etchemendy, Ligett, Lyons, Manyika, Niebles, Shoham, Wald, and Clark, 2024), industry participation in responsible AI research remains comparatively limited. This pattern suggests that current incentive structures may not fully align with the development of responsible AI (Askell, Brundage, and Hadfield, 2019). Prior research suggests that competitive pressures may accelerate the commercialization process, potentially affecting investment in responsible AI development (Askell, Brundage, and Hadfield, 2019). Regulatory frameworks represent one mechanism that could encourage AI firms to demonstrate engagement in responsible AI research as part of product development and deployment processes.
@@
Transparency and public accountability represent important dimensions of the AI development process (Brundage, Avin, Wang, Belfield, Krueger, Hadfield, Khlaaf, Yang, Toner, Fong, Maharaj, Koh, Hooker, Leung, Trask, Bluemke, Lebensold, O’Keefe, Koren, and Ryffel, 2020; Mittelstadt, 2019). To foster trust and confidence in the AI ecosystem, policymakers may consider mechanisms for AI firms to disclose their responsible AI research efforts, ethical frameworks, and decision-making processes related to the development and deployment of AI systems (de Laat, 2021; de Almeida, dos Santos, and Farias, 2021). Such disclosures could be complemented by independent audits and public review processes, supporting alignment between AI development and societal values and interests (Mökander, Morley, Taddeo, and Floridi, 2021).
@@
Furthermore, allocation of funding and resources toward independent, public-sector responsible AI research initiatives represents another policy consideration. While industry participation contributes to the field, our findings highlight the value of maintaining a robust research ecosystem that can examine AI technologies and their societal impacts from perspectives independent of commercial considerations (Baker and Hanna, 2022).
@@
Industry-academia research collaborations in responsible AI represent one mechanism that could help bridge observed gaps. Such partnerships may promote knowledge exchange, foster the integration of academic insights into industry practices, and support a culture of responsible AI development within firms. By leveraging the combined expertise and resources of academia and industry, such collaborations could contribute to the advancement of responsible AI research and its practical implementation.
7.2 Implications for the Direction of AI Development